OpenStudy (anonymous):

Please help F(x) = -(x-1)2+6 F(x) = 4(x+1) - 1 Solve for x : 3(1+2x) = 27

OpenStudy (anonymous):

Do they want you to solve the last equation there? In order to do that we don't need the first two functions F you have listed. Is this how the question is laid out?

OpenStudy (anonymous):

Need to solve 3 equations

OpenStudy (anonymous):

No

OpenStudy (anonymous):

1) F(x) = -(x-1)2+6 2) F(x) = 4(x+1) - 1 3) Solve for x : 3(1+2x) = 27

OpenStudy (anonymous):

Ok for 3) we follow the following steps: 3+6x=27 6x=24 x=4 For equations 1) and 2), are you sure they don't give you what F(x) is meant to equal? It's not possible to just "solve" functions that are given that way.

OpenStudy (anonymous):

I'm sure

OpenStudy (anonymous):

Thank you for #3, I was correct.

OpenStudy (anonymous):

actually on #1 the 2 is an Exponent and on # 2 the (x+4) is an exponent